### v1

# A Generalized Nonlinear extension of Quantum Mechanics

Abstract
We construct the most general form of our previously proposed nonlinear extension of quantum mechanics that possesses three basic properties. Unlike the simpler model, the new version is not completely integrable, but it has an underlying Hamiltonian structure. We analyze a particular solution in detail, and we use a natural extension of the Born rule to compute particle trajectories. We find that closed particle orbits are possible.
pacs:
3.65Ud, 3.65Ta, 05.45.-a, 11.10.Lm
I Introduction of the model
In a previous paper cf23 , we proposed a non-linear extension of quantum mechanics and studied some of its consequences. Our motivations were various: partly, in the spirit of Weinberg Weinberg89 and Kaplan and Rajendran Kaplan , we simply wanted to see what a particular non-linear extension of quantum mechanics might look like, and to deduce some of its consequences.
But we also had in mind the question of whether, as others have speculated Penrose Carlip , the union of gravity with quantum mechanics is best achieved not by producing a fully quantum version of general relativity, as is contemplated for example in string theory or loop quantum gravity, but by some other means, such as the coupling of quantum mechanics to a classical version of gravity, as in the work of Oppenheim and collaborators Oppenheim , or indeed by the modification of quantum mechanics itself.
In the model we introduced in cf23 ; cf24 , a given system is described by 2 state vectors, $|\psi>$ and $|\phi>$, in Hilbert space, in contrast to the single state vector to which ordinary quantum mechanics limits itself. In cf23 an exploration of the extra information provided by the pair of state vectors was not attempted. The model had the advantage of simplicity and was completely integrable, in the sense that the equations of motion determined the time dependence of the inner products $\psi|\psi>$, $\phi|\psi>$, and $\phi|\phi>$ explicitly in terms of a small number of integration constants, and one could then go on to determine the modification of the time dependence of the state vectors themselves due to the non-linearities. The non-linearity in that model is proportional to the inner product $\gamma=<\phi|\psi>$. One finds that $\gamma\rightarrow 0$ for large times, so that the effect of the non-linearity vanishes in the far future and the far past. In this work we seek a generalization that will allow for a broader range of dynamical possibilities. If we define
$|\Psi>=\begin{pmatrix}|\psi>\\ |\phi>\\ \end{pmatrix}$ (1)
and
$\mathcal{H=}\begin{pmatrix}H&0\\ 0&H\\ \end{pmatrix}$ (2)
where $H$ is the Hamiltonian of the underlying system, then the equations of motion of the model in reference cf23 are
$i\frac{\partial|\Psi>}{\partial t}\mathcal{=H|}\Psi\mathcal{>+M|}\Psi>$ (3)
where $\mathcal{M}$ is the Hermitian matrix
$\mathcal{M=}\begin{pmatrix}0&g<\phi|\psi>\\ g^{*}<\psi|\phi>&0\\ \end{pmatrix}$ (4)
The coupling constant $g=a+ib$ is in general complex ($b\neq 0)$ .
[…]
  1. (I)
    Each element of $\mathcal{M}$ should depend linearly on the inner products $\psi|\psi>$, $\phi|\psi>$, $\psi|\phi>$, and $\phi|\phi>$, and should depend on time only through them;
  2. (II)
    The equations of motion should imply that $\psi|\psi>+<\phi|\phi>=N=const.;$
  3. (III)
    The equations of motion should be invariant under the separate constant phase transformations $\left|\psi>\rightarrow e^{i\theta_{1}}\right|\psi>$ and $\left|\phi>\rightarrow e^{i\theta_{2}}\right|\phi>$ .
To implement condition (II), it is sufficient for $\mathcal{M}$ to be Hermitian, but it is not necessary. $\mathcal{M}$ can have an anti-Hermitian part of the form
$\mathcal{M}^{{}^{\prime}}=\begin{pmatrix}i\mu<\phi|\phi>&\lambda<\phi|\psi>\\ -\lambda<\psi|\phi>&-i\mu<\psi|\psi>\\ \end{pmatrix}$, (5)
where $\mu$ and $\lambda$ are real constants. Then the most general $\mathcal{M}$ satisfying (I), (II), and (III) is
$\mathcal{M=}\mathcal{M}_{0}\mathcal{+M^{\prime}}$ (6)
where
$\mathcal{M}_{0}=\mathcal{M}_{0}^{\dagger}=\begin{pmatrix}\alpha_{1}<\psi|\psi>+\alpha_{2}<\phi|\phi>&g<\phi|\psi>\\ g^{*}<\psi|\phi>&\beta_{1}<\psi|\psi>+\beta_{2}<\phi|\phi>\\ \end{pmatrix}$ (7)
Here the $\alpha_{i}$ and $\beta_{i}$ are real constants, and $g=a+ib$ was defined above. In total, there are 8 real parameters, as compared to two in the original model.
I.1 The $\tau$- $\delta$ problem
If $\mathcal{M}=0$, all the inner products would be constant. Hence, the inner products obey evolution equations that are independent of $H.$ In these equations, we find that most of the above parameters play no role. With $\gamma=<\phi|\psi>,\$we define $\delta=|\gamma|^{2}$, and let
$\tau=<\psi|\psi>-<\phi|\phi>.$ (8)
We then have that $\tau$ and $\delta$ obey the following differential equations:
$\frac{d\tau}{dt}$ $=\mu\left(N^{2}-\tau^{2}\right)+4b\delta,$
$\frac{d\delta}{dt}$ $=-2(b+\mu)\tau\delta.$ (9)
The analogous equations for the original model are recovered for $\mu=0$. To gain some preliminary insight, we note that we can easily integrate eqs.(I.1) in 3 special cases: (1) $\mu=0$;   (2) $b+\mu=0;\$  (3) $b=0$ .
Case (1) is just the original model, for which we have that $\tau(t)$ and $\delta(t)$ are given by:
[…]
where $\omega_{0}$ is a constant of integration. In general, we expect the solution of these 2 equations to depend on one constant of integration in addition to the freedom to redefine the origin of time, $t\rightarrow t-t_{0},$ which we do not explicitly indicate.
[…]
Noting that in all three special cases $\tau$ is a hyperbolic tangent, we can make the ansatz
[…]
where $\alpha^{\prime}$ and $\beta^{\prime}\$are constants to be determined, and plug this into the full equations for $\tau$ and $\delta$. We indeed find a solution, namely
[…]
and
[…]
But this is definitely not the general solution, because $\alpha^{\prime}$ and $\beta^{\prime}$ are fixed by the equations, so there is no free constant of integration. Indeed, if we go through the special cases that we solved above, we find that our solution to the general equations corresponds to a specific choice for the constant of integration: Case (1), $\omega_{0}=\frac{N}{2}$; Case (2), $\omega_{0}=0$; Case (3), $\delta_{0}=\frac{N^{2}}{4}.$
II Hamiltonian formulation of the $\tau$ -$\delta$ system of equations
[…]
$\frac{d\kappa}{dt}=-2(b+\mu)\tau$ (14)
and therefore
$\frac{d\tau}{dt}=\frac{d\tau}{d\kappa}\frac{d\kappa}{dt}=-(b+\mu)\frac{d\Omega}{d\kappa}=\mu\left(N^{2}-\Omega\right)+4be^{\kappa}.$ (15)
This is a first-order linear differential equation for $\Omega$ as a function of $\kappa$, and can be integrated exactly:
$\Omega=\tau^{2}=N^{2}-4e^{\kappa}-ce^{\frac{\mu\kappa}{b+\mu}}=N^{2}-4\delta-c\delta^{\frac{\mu}{b+\mu}}.$ (16)
Here $c$ is a constant of integration. In the original model, $\mu=0,$ and $c=N^{2}-4\omega_{0}^{2}$ . The relation above is the generalization of $\omega_{0}^{2}=\frac{\tau^{2}}{4}+\delta$ (eq. (2.15) of cf23 ) in the original model. Since
$N^{2}-\tau^{2}-4\delta=4\mathcal{S},$ (17)
where
$\mathcal{S}=<\psi|\psi><\phi|\phi>-<\psi|\phi><\phi|\psi>$ (18)
is the Schwarz parameter, the relation can also be expressed as
$4\mathcal{S}=c\delta^{\frac{\mu}{b+\mu}}.$
In the original model, $\mathcal{S}$ is a constant, but in this more general case it apparently will not be.
To solve these equations completely, we would need to substitute the relation between $\tau$ and $\delta$ back into the equations for $\frac{d\tau}{dt}$ and $\frac{d\delta}{dt}$ and perform one more integration.
To simplify the subsequent discussion, we adopt a slightly different notation. As evolution parameter, we choose $s=-(b+\mu)t$, and define $p=\frac{\mu}{(b+\mu)}$.[Note: this requires $b+\mu\neq 0,\$so Case 2 above is not covered by this analysis.] The equations of motion for $\tau$ and $\kappa$ are then
$\frac{d\kappa}{ds}=2\tau$ (19)
and
$\frac{d\tau}{ds}=-4e^{\kappa}-pce^{p\kappa},$ (20)
with the additional relation
$N^{2}=h\equiv\tau^{2}+4e^{\kappa}+ce^{p\kappa}.$ (21)
We recognize that $h$ serves as the Hamiltonian for this system, with $\kappa$ playing the role of the coordinate and $\tau$ the role of the momentum. In fact the first order equations for $\kappa$ and $\tau$ are just Hamilton’s equations:
$\frac{d\kappa}{ds}=\frac{\partial h}{\partial\tau}$ (22)
and
$\frac{d\tau}{ds}=-\frac{\partial h}{\partial\kappa}$ (23)
So $\kappa$ behaves like the position $x$ of a particle in 1-D, moving in the potential $V(\kappa)=4e^{\kappa}+ce^{p\kappa}$. This explains in a more general way why in the original model $\delta=|\gamma|^{2}\rightarrow 0$ for large times. In the original model, $p=0,$ so $V(\kappa)$ tends to infinity for large positive $\kappa$, but for large negative $\kappa$ the potential tends smoothly to a constant. Therefore $\kappa$ escapes to minus infinity, and $\delta=e^{\kappa}$ tends to zero.
If $p>0\ ,\ \ \ V(\kappa)=4e^{\kappa}+ce^{p\kappa}$ will have the same general behavior as in the $p=0$ case. (The parameter $c$ is constrained to be positive because of the relation $ce^{p\kappa}=4\mathcal{S).}$ But if $p<0$, $V(\kappa)$ will tend to infinity both for large positive and negative $\kappa$, insuring that $\delta$ stays bounded away from either zero or infinity as time evolves. In terms of the original parameters, assuming $b>0$, this means $b<\mu<0$ . This is the range of parameters we shall consider going forward.
It is not known to us whether the problem of a classical 1-D particle moving in a potential $V(\kappa)=4e^{\kappa}+ce^{p\kappa}$ with c>0c>0\ and $p<0$ can be solved exactly. Below we present the general solution for a particular choice of parameters, and another particular solution that is valid for all choices of the parameters.
Energy conservation leads to
$E=N^{2}=\tau^{2}+4e^{\kappa}+ce^{p\kappa},$ (24)
where from Hamilton’s equations:
$2\tau=\frac{d\kappa}{ds}.$ (25)
Thus
$N^{2}=\frac{1}{4}(\frac{d\kappa}{ds})^{2}+V(\kappa)$ (26)
and
$ds=\pm\frac{d\kappa}{2\sqrt{N^{2}-V(\kappa)}}.$ (27)
II.1 The case p=1p=-1
If we choose $\mu=-\frac{b}{2}$, then p=1p=-1. This also leads to $b+\mu=\frac{b}{2}$. so that $s=-\frac{b}{2}t$ Furthermore, we can translate the coordinate: $\kappa=\eta+\eta_{0}$, where $\eta_{0}=\frac{1}{2}\ln\frac{c}{4}$ , so that the potential becomes $V(\eta)=4{\sqrt{c}\ \cosh}\eta$.
The potential problem we want to solve is
$\frac{d^{2}\eta}{ds^{2}}=-\frac{dV}{d\eta}=-4\sqrt{c}\sinh(\eta).$ (28)
The solution is
$\eta=\pm 2i~\text{am}\left(\frac{1}{2}\sqrt{\left(8\sqrt{c}-c_{1}\right)(s+c_{2}){}^{2}}|-\frac{16\sqrt{c}}{c_{1}-8\sqrt{c}}\right)$ (29)
Here am $(u|m)$ is the Jacobi amplitude which is the inverse of the incomplete elliptic integral of the first kind $F(u|m)$. Note that to obtain this symmetric potential we needed to choose $\mu=-b/2$ so that the parameter $s=-\frac{bt}{2}$ where $t$ is the actual time.
Utilizing Energy conservation Eq. (24) we determine these constants. In particular one has for the ++ sign in Eq. (27)
$s=-\frac{i~F\left(\frac{i\kappa}{2}|\frac{8\sqrt{c}}{4\sqrt{c}-N^{2}}\right)}{\sqrt{N^{2}-4.\sqrt{c}}}$ (30)
Inverting this equation one obtains:
$\kappa=-2i~\text{am}\left(i\sqrt{N^{2}-4\sqrt{c}}s|\frac{8\sqrt{c}}{4\sqrt{c}-N^{2}}\right)$ (31)
which is shown in Fig.( 1) for $c=4,N^{2}=25$
Figure 1: $\kappa(s)$ for $N^{2}=25,c=4$
From $\kappa$ we can determine $\tau$
$\tau(s)=\frac{1}{2}\frac{d\kappa}{ds}=\sqrt{N^{2}-4\sqrt{c}}~\text{dn}\left(i\sqrt{N^{2}-4\sqrt{c}}s|\frac{8\sqrt{c}}{4\sqrt{c}-N^{2}}\right)$ (32)
where $\text{dn}(u|m)$ is the delta amplitude, satisfying the relationship
$\text{dn}(u|m)=\frac{d}{du}\text{am}(u|m).$
When $N^{2}=25.c=4$ we get the behavior for $\tau(s)$ shown in Fig. (2) .
Figure 2: $\tau(s)$ for $N^{2}=25,c=4$
For the same parameters we find that $\delta(s)=e^{\kappa}$ has the behavior shown in Fig. (3).
Figure 3: $\delta(s)$ for $N^{2}=25,c=4$
Here, $s=-(b+\mu)t=-\frac{1}{2}bt$ and $N^{2}=25,c=4$.
II.2 A simple solution when the “momentum” $\tau$ is zero
Given the shape of the potential, there is one obvious, simple, and surprisingly non-trivial solution, when the particle sits at the bottom of the potential with zero momentum $\tau$.
From the equations above, the conditions for this solution are:
$\frac{d\tau}{ds}=-4e^{\kappa}-pce^{p\kappa}=0$ (33)
and $N^{2}=4e^{\kappa}+ce^{p\kappa}$. These two conditions determine $c$ and $e^{\kappa}=\delta$. In particular, we find from Eq. (15) and the fact $\Omega=0$, that
$\delta=\left(\frac{-\mu}{4b}\right)N^{2}.$ (34)
(Reminder: $\mu<0,$ so $\delta>0$ as it must be.) Also, since $\tau=0$, we have
$\psi|\psi>=<\phi|\phi>=\frac{N}{2}.$
From the original equations of motion, we find that
$i\frac{d\gamma}{dt}=\left[\lambda N-(g+i\mu)\tau+\frac{1}{2}\left(\alpha_{1}-\beta_{1}\right)(N+\tau)+\frac{1}{2}\left(\alpha_{2}-\beta_{2}\right)(N-\tau)\right]\gamma,$ (35)
which in our case reduces to
$i\frac{d\gamma}{dt}=\frac{N}{2}[2\lambda+\alpha-\beta]\gamma,$ (36)
where we have abbreviated $\alpha_{1}+\alpha_{2}=\alpha$ and $\beta_{1}+\beta_{2}=\beta$. Defining $\theta=\frac{N}{2}[2\lambda+\alpha-\beta]$ we have
$\gamma=\gamma_{0}e^{-i\theta t}\,$ (37)
with ${\gamma_{0}}^{2}=\ \delta=\left(\frac{-\mu}{4b}\right)N^{2}$ . (In principle, $\gamma$ also contains an overall constant phase, which we have here neglected.)
With $\gamma$ in hand, we proceed to the equations of motion for the state vectors, which in the current case become:
$i\frac{d|\psi>}{dt}=H|\psi>+\frac{N}{2}(\alpha+i\mu)|\psi>+(g+\lambda)\gamma|\phi>$ (38)
and
$i\frac{d|\phi>}{dt}=H|\phi>+\frac{N}{2}(\beta-i\mu)|\phi>+\left(g^{*}-\lambda\right)\gamma^{*}|\psi>.$ (39)
It is convenient to expand both state vectors in eigenstates of $H$ (assuming a discrete spectrum for notational convenience):
$|\psi>=\sum_{n}{\psi_{n}e^{-i{(E}_{n}+q)t}}|n>$ (40)
and
$|\phi>=\sum_{n}{\phi_{n}e^{-i{(E}_{n}+q^{{}^{\prime}})t}}|n>$ (41)
where $q=\frac{N}{2}(\alpha+i\mu)$ and $q^{{}^{\prime}}=\frac{N}{2}(\beta-i\mu)$ . We expect that the imaginary parts of $q$ and $q^{{}^{\prime}}$ will cancel in the final expressions for $|\psi>$ and $|\phi>$.
We have
$i{\dot{\psi}}_{n}=(g+\lambda)\gamma_{0}e^{-i\theta^{{}^{\prime}}t}\phi_{n}$ (42)
and
$i{\dot{\phi}}_{n}=\left(g^{*}-\lambda\right)\gamma_{0}e^{i\theta^{{}^{\prime}}t}\psi_{n},$ (43)
with $\theta^{{}^{\prime}}=\ N(\lambda-i\mu)$.
We can decouple these equations by taking another time derivative:
${\ddot{\psi}}_{n}+iN(\lambda-i\mu){\dot{\psi}}_{n}+(g+\lambda)\left(g^{*}-\lambda\right){\gamma_{0}}^{2}\psi_{n}=0.$ (44)
This has solutions of the form $\psi_{n}=K_{n}e^{\nu t}$, where $\nu$ satisfies
$\nu^{2}+i\ N(\lambda-i\mu)\nu+(g+\lambda)\left(g^{*}-\lambda\right){\gamma_{0}}^{2}=0.$ (45)
The two solutions for $\nu$ are:
$\nu_{\pm}=-\frac{\mu N}{2}+i\chi_{\pm},$ (46)
where $\chi_{\pm}=-\frac{\lambda N}{2}\pm\frac{N}{2}\left|\sqrt{{\frac{\mu}{b}a^{2}+\mu}^{2}\left(1+\frac{b}{\mu}\right)-\lambda^{2}\left(1+\frac{\mu}{b}\right)}\ \right|\equiv-\frac{\lambda N}{2}\pm\sigma$.
With our choice of parameters, $b<\mu<0$, all 3 terms under the radical are negative. Therefore (we will need this later)
$\sigma^{2}=\frac{N^{2}}{4}\left(-\frac{\mu}{b}a^{2}-\mu^{2}\left(1+\frac{b}{\mu}\right)+\lambda^{2}\left(1+\frac{\mu}{b}\right)\right).$ (47)
With these inputs, we find that
$|\psi>=\sum_{n}e^{-i\left(E_{n}+\frac{N}{2}\alpha+\lambda\frac{N}{2}\right)t}\left[K_{n}^{(+)}e^{i\sigma t}+K_{n}^{(-)}e^{-i\sigma t}\right]|n>$ (48)
and
$|\phi>=\sum_{n}e^{-i\left(E_{n}+\frac{N}{2}\beta-\lambda\frac{N}{2}\right)t}\left[K_{n}^{{}^{\prime}(+)}e^{i\sigma t}+K_{n}^{{}^{\prime}(-)}e^{-i\sigma t}\right]|n>.$ (49)
To solve the equations of motion, we must impose the following relation between the $K_{n}$ and the $K_{n}^{{}^{\prime}}$:
$K_{n}^{{}^{\prime}(\pm)}=\frac{i\nu_{\pm}}{(g+\lambda)\gamma_{0}}K_{n}^{(\pm)}.$ (50)
Note that this proportionality factor does not depend on $n$.
The $K_{n}^{(\pm)}$ are not completely arbitrary, because we must insure that
$\psi|\psi>=<\phi|\phi>=\frac{N}{2}$ (51)
and
$\phi|\psi>=\gamma_{0}e^{-i\theta t},$ (52)
where $\theta=\frac{N}{2}[2\lambda+\alpha-\beta]$. .As we shall see, these are 4 real conditions for only 2 variables, so it needs to be checked that they can indeed be satisfied.
In evaluating these conditions, there will be time-dependent cross terms that can be eliminated by imposing
$\sum_{n}{K_{n}^{(-)*}K_{n}^{(+)}}=\sum_{n}{K_{n}^{(+)*}K_{n}^{(-)}=0}.$ (53)
This also implies
$\sum_{n}{K_{n}^{{}^{\prime}(-)*}K_{n}^{{}^{\prime}(+)}}=\sum_{n}{K_{n}^{{}^{\prime}(+)*}K_{n}^{{}^{\prime}(-)}=0}.$ (54)
Define $S_{\pm}=\sum_{n}{K_{n}^{(\pm)*}K_{n}^{(\pm)}}$. Then we have the conditions
$\psi|\psi>=\frac{N}{2}\ \ \Longrightarrow\ \ \ \ \ \ S_{+}+S_{-}=\frac{N}{2},$ (55)
$\phi|\phi>=\frac{N}{2}\ \ \Longrightarrow\ \ \ \ \ \ \ \left|\nu_{+}\right|^{2}{\ S}_{+}+\left|\nu_{-}\right|^{2}\ S_{-}\ \ =\ \frac{N}{2}\ (g^{*}+\lambda)(g+\lambda){\gamma_{0}}^{2},$ (56)
and
$\phi|\psi>=\gamma\ \Longrightarrow\ \ \ \ \ \ \ \nu_{+}^{*}S_{+}+\nu_{-}^{*}S_{-}=i{\gamma_{0}}^{2}(g^{*}+\lambda).$ (57)
These are 4 real equations for the 2 unknowns, $S_{+}$ and $S_{-}$. We can dispose of one of these equations immediately. The real parts of $\nu_{+}^{*}$ and $\nu_{-}^{*}$ are both $\frac{\mu N}{2}$, and the real part of the right-hand side of Eq. (57) (using ${\gamma_{0}}^{2}=\left(\frac{-\mu}{4b}\right)N^{2}$) is $\left(\frac{-\mu}{4}\right)N^{2}$, so the real part of that equation is obeyed.
The imaginary part of that equation yields an expression for $S_{+}-S_{-}$ :
$S_{+}-S_{-}=\frac{N^{2}}{4\sigma}\left[\lambda\left(1+\frac{\mu}{b}\right)+a\frac{\mu}{b}\right].$ (58)
We can obtain another expression for $S_{+}-S_{-}$ from Eq. (56) :
$S_{+}-S_{-}=\frac{1}{4\lambda\sigma}\left[\frac{N^{2}}{2}\left(\mu^{2}+\lambda^{2}\right)+2\sigma^{2}+\frac{N^{2}}{2}\frac{\mu}{b}\left(a^{2}+b^{2}+2\lambda a+\lambda^{2}\right)\right].$ (59)
Using the expression above for $\sigma^{2}$ and some algebra, one can show that these are the same, allowing for a consistent solution for $|\psi>$ and $|\phi>.$
One remaining consistency check is that, by definition, $S_{+}$ and $S_{-}$ must both be positive, which requires $\left|S_{+}-S_{-}\right|<\frac{N}{2}$. From Eq.( 58) , this amounts to
$\frac{4}{N^{2}}\sigma^{2}-\left[\lambda\left(1+\frac{\mu}{b}\right)+a\frac{\mu}{b}\right]^{2}>0.$ (60)
Working out the expression on the left-hand side using the formula for $\sigma^{2}$ given above, we find that it is equal to
$\frac{\mu}{b}\left(1+\frac{\mu}{b}\right)(a+\lambda)^{2}-\mu(\mu+b),$ (61)
which is manifestly positive because $b<\mu<0$ .
We can define two vectors that evolve according to $H$, independently of the non-linear terms:
$|A>=\sum_{n}{K_{n}^{(+)}e^{-iE_{n}t}}|n>$ (62)
and
$|B>=\sum_{n}{K_{n}^{(-)}e^{-iE_{n}t}}|n>,$ (63)
and then write
$|\psi>=e^{-i\frac{N}{2}(\alpha+\lambda)t}\left\{e^{i\sigma t}|A>+e^{-i\sigma t}|B>\right\}$ (64)
and
$|\phi>=\frac{i}{(g+\lambda)\gamma_{0}}e^{-i\frac{N}{2}(\beta-\lambda)t}\left\{\nu_{+}e^{i\sigma t}|A>+\nu_{-}e^{-i\sigma t}|B>\right\}.$ (65)
$|A>$ and $|B>$ must obey the normalization and orthogonality conditions
$A|A>+<B|B>=\frac{N}{2}\ \ \ \ \ ;\ \ \ \ <A|A>-<B|B>\ =\ \frac{N^{2}}{4\sigma}\left[\lambda\left(1+\frac{\mu}{b}\right)+a\frac{\mu}{b}\right]\$(66)
and <A|B>=0.<A|B>=0\ .
These equations express the general solution for $|\psi>$ and $|\phi>$, for the particular solution of the $(\tau,\delta)$ problem that we have considered ($\tau=0,\ \delta=const.$). The fact that $|\psi>$ is expressed more simply than $|\phi>$ is an artifact of the order in which we solved the equations and is not a basic property of the system.
III Properties of the density matrix and the trajectory function
In cf23 we defined the density matrix
$\rho\ =\frac{1}{N}\left[|\psi><\psi|+|\phi><\phi|\right],$ (67)
and observed that it obeys the equation $i\dot{\rho}=[H,\rho]$, i.e. all the non-linear effects disappear from $\rho$. This raises a problem, because the most natural extension of the Born rule to our model is to use $\rho$ to define the probability. However, the effects of the non-linearity would not then be manifest in physical quantities.
Is the same true in the extended model? Using the full equations of motion for $|\psi>$ and $|\phi>$ , we find
$~~i\dot{\rho}$ $=[H,\rho]+2i\mu\left[\left|\psi><\phi|\phi><\psi\right|-|\phi><\psi|\psi><\phi|\right]$ (68)
$+2\lambda\left[\left|\phi><\phi|\psi><\psi\right|-|\psi><\psi|\phi><\phi|\right].$
In the case at hand, this becomes
$i\dot{\rho}=[H,\rho]+i\mu N\left[|\psi><\psi|-|\phi><\phi|\right]+2\lambda\left[\gamma|\phi><\psi|-\gamma^{*}|\psi><\phi|\right].$ (69)
We also have
$Tr\ \rho^{2}=1-\frac{2\mathcal{S}}{N^{2}}=1-\frac{c}{{2N}^{2}}\delta^{\frac{\mu}{b+\mu}}.$ (70)
Since $\rho$ now includes effects of the non-linearity, it makes sense to define the expectation value of an operator $X$ as
$X(t)>=Tr\ \rho(t)X=\frac{1}{N}\left[<\psi(t)|X\left|\psi(t)>+<\phi(t)\right|X|\phi(t)>\right].$ (71)
In what follows, we shall be especially interested in the case where <X(t)><X(t)> represents a trajectory, i.e. where $X$ is the position operator for some system. In a somewhat different context, we pursued the idea of using trajectories of this type to investigate the classical effects of the non-linearity in cf24 .
To isolate the effects of the non-linearity, we want the system to move as an otherwise free particle. We therefore take the Hamiltonian $H$ to be a function of $P$ only, where $P$ is the momentum conjugate to $X$. It follows that for any two vectors $|\psi_{1}>$ and $|\psi_{2}>$ that evolve according to
$\left|\psi_{i}(t)>\ =e^{-iHt}\right|\psi_{i}>,\ \ \ \ \ i=1,2$ (72)
we have
$\frac{d^{2}}{dt^{2}}<\psi_{2}|X|\psi_{1}>=0$ (73)
and therefore
$\psi_{2}|X|\psi_{1}>=c_{1}t+c_{2}.$ (74)
If ${|\psi}_{1}>=|\psi_{2}>$, $c_{1}$ and $c_{2}$ must be real constants. Otherwise, they can be complex.
III.1 Trajectory function for the simple solution
We would like to calculate the trajectory for the case of the “simple solution” we have obtained above. We consider a position operator in 3 dimensions, $X_{i}$, $i=1,2,3.$ The quantities $A\left|X_{i}\right|A>,\ <A\left|X_{i}\right|B>$ and $B\left|X_{i}\right|B>$ will all be linear functions of time. We are interested in the closed orbit of the particle, not in the overall motion of the system through space (the analogy might be that we are interested in the orbit of Earth around the sun, not the motion of the solar system through space). Hence, we take these matrix elements to be constants.
Using the explicit formulas for $|\psi>$ and $|\phi>$ , we can express $\overrightarrow{X}(t)>$ as
[…]
where ${\overrightarrow{X}}_{0}$, $\overrightarrow{V}$ and $\overrightarrow{W}$ are constant vectors depending on the values we choose for $A\left|X_{i}\right|A>,\ <A\left|X_{i}\right|B>$ and $B\left|X_{i}\right|B>$ . Thus, the particle moves in the plane spanned by $\overrightarrow{V}$ and $\overrightarrow{W}$. Calling the coordinates in this plane $X_{1}^{{}^{\prime}}$ and $X_{2}^{{}^{\prime}}$, we can orient the axes so that the equation for the orbit can be written in the form
$\frac{{X_{1}^{{}^{\prime}}}^{\mathbf{2}}}{{R_{1}}^{\mathbf{2}}}\mathbf{+}\frac{{X_{2}^{{}^{\prime}}}^{\mathbf{2}}}{{R_{2}}^{\mathbf{2}}}\mathbf{=}1,$ (76)
[…]
We expect, however, that for more general solutions to the tau-delta problem, the trajectory motion will be less symmetrical, because the potential itself is asymmetric (except for p=1p=-1).
III.2 General formalism
Following the same strategy as for the simple solution where $\kappa$ is a constant, in this section we derive the equations that will be satisfied in the general case, although of course we do not (yet) have solutions for these equations.
The first step is to solve the “tau-delta” problem, i.e. the pair of coupled, non-linear equations
$\frac{d\kappa}{ds}=2\tau$ (77)
and
$\frac{d\tau}{ds}=-4e^{\kappa}-pce^{p\kappa},$ (78)
together with the constraint
$N^{2}=h\equiv\tau^{2}+4e^{\kappa}+ce^{p\kappa}.$ (79)
Here $s=-(b+\mu)t$ and $h$ is the Hamiltonian:
$\frac{d\kappa}{ds}=\frac{\partial h}{\partial\tau}$ (80)
and
$\frac{d\tau}{ds}=-\frac{\partial h}{\partial\kappa}.$ (81)
The equation for $\phi|\psi>=\gamma$ is
$i\frac{d\gamma}{dt}=\left[\lambda N-(g+i\mu)\tau+\frac{1}{2}\left(\alpha_{1}-\beta_{1}\right)(N+\tau)+\frac{1}{2}\left(\alpha_{2}-\beta_{2}\right)(N-\tau)\right]\gamma,$ (82)
which can be solved exactly. We define $\alpha_{1}+\alpha_{2}=\alpha$, $\beta_{1}+\beta_{2}=\beta$, $\alpha_{1}-\alpha_{2}=\alpha^{\prime},$ and $\beta_{1}-\beta_{2}=\beta^{\prime}$. The solution is
$\gamma=e^{\frac{\kappa}{2}}e^{-i(\Lambda+\theta_{0})},$ (83)
where $\Lambda=N\left[\lambda+\frac{1}{2}(\alpha-\beta)\right]t+\frac{\kappa}{2(b+\mu)}\left[a-\frac{1}{2}(\alpha^{\prime}-\beta^{\prime})\right]$, and $\theta_{0}$ is an arbitrary real constant. We shall suppress $\theta_{0}$ in the following analysis. The equations for the state vectors are
$i\frac{d|\psi>}{dt}=\left[H+\frac{N}{2\ }((\alpha+i\mu)+\frac{\tau}{2}(\alpha^{\prime}-i\mu)\right]\left|\psi>+(g+\lambda)\gamma\right|\phi>$ (84)
and
$i\frac{d|\phi>}{dt}=\left[H+\frac{N}{2\ }((\beta-i\mu)+\frac{\tau}{2}(\beta^{\prime}-i\mu)\right]\left|\phi>+\left(g^{*}-\lambda\right)\gamma^{*}\right|\psi>.$ (85)
We define
$|\psi>=e^{\frac{\mu Nt}{2}+\frac{\mu\kappa}{4(b+\mu)}}e^{-i\theta_{\psi}}|\psi^{{}^{\prime}}>,$ (86)
where $\theta_{\psi}=\left[H+\frac{N}{2}\alpha\right]t-\frac{\kappa\alpha^{{}^{\prime}}}{4(b+\mu)}$ , and also
$|\phi>=e^{\frac{-\mu Nt}{2}+\frac{\mu\kappa}{4(b+\mu)}}e^{-i\theta_{\phi}}{|\phi}^{{}^{\prime}}>,$ (87)
where $\theta_{\phi}=\left[H+\frac{N}{2}\beta\right]t-\frac{\kappa\beta^{{}^{\prime}}}{4(b+\mu)}$. Then we find the coupled equations
$i\frac{d|\psi^{\prime}>}{dt}=e^{f}(g+\lambda)|\phi^{{}^{\prime}}>$ (88)
and
$i\frac{d|\phi^{\prime}>}{dt}=e^{\kappa-f}\left(g^{*}-\lambda\right)|\psi^{{}^{\prime}}>,$ (89)
[…]
We can decouple these by taking another derivative:
$\frac{d^{2}|\psi^{{}^{\prime}}>}{dt^{2}}-\frac{df}{dt}\frac{d|\psi^{\prime}>}{dt}+(g+\lambda)\left(g^{*}-\lambda\right)e^{\kappa}|\psi^{{}^{\prime}}>=0.$ (90)
There is a similar equation for $\frac{d^{2}|\phi^{{}^{\prime}}>}{dt^{2}}$ , in which $f$ is replaced by $f^{{}^{\prime}}=\kappa-f$. When $\kappa=const.,$ we find agreement with Eq. ( 44) above. If we assume a solution of the form
$|\psi^{{}^{\prime}}>=F(t)|\psi_{0}>$ (91)
where $|\psi_{0}>$ is a constant vector, then $F(t)$ obeys
$\frac{d^{2}F}{dt^{2}}-\frac{df}{dt}\frac{dF}{dt}+(g+\lambda)\left(g^{*}-\lambda\right)e^{\kappa}F=0.$ (92)
This is a second-order linear differential equation, so it has 2 linearly independent solutions, $F_{1}$ and $F_{2}$. The general solution for $|\psi^{{}^{\prime}}>$ is then
$|\psi^{{}^{\prime}}>=F_{1}(t)|\psi_{1}>+F_{2}(t)|\psi_{2}>,$ (93)
where the $|\psi_{i}>$ are arbitrary constant vectors. It follows that
$|\phi^{\prime}>=\frac{e^{-f}}{(g+\lambda)}i\left[\frac{dF_{1}}{dt}|\psi_{1}>+\frac{dF_{2}}{dt}|\psi_{2}>\right].$ (94)
The Wronskian of $F_{1}$ and $F_{2}$ has a simple form:
$W\left(F_{1},F_{2}\right)\equiv\dot{F_{1}}F_{2}-F_{1}\dot{F_{2}}=c_{0}e^{f},$ (95)
where $c_{0}$ is a constant of integration. To recover $|\psi>$ and $|\phi>$ we first define
$|\psi_{i}(t)>=e^{-iHt}|\psi_{i}>\ ,\ \ i=1,2.$ (96)
These are 2 vectors that evolve according to the original Hamiltonian $H$. We then have
$|\ \psi>=e^{-i\frac{N}{2}(\alpha+i\mu)t}e^{\frac{i\kappa}{4}\frac{(\alpha^{{}^{\prime}}-i\mu)}{(b+\mu)}}\left[F_{1}(t)|\psi_{1}(t)>+F_{2}(t)|\psi_{2}(t)>\right]$ (97)
and
$|\ \phi>=e^{-i\frac{N}{2}(\beta-i\mu)t}e^{\frac{i\kappa}{4}\frac{(\beta^{{}^{\prime}}-i\mu)}{(b+\mu)}}\frac{e^{-f}}{(g+\lambda)}i\left[\frac{dF_{1}}{dt}|\psi_{1}(t)>+\frac{dF_{2}}{dt}|\psi_{2}(t)>\right].$ (98)
If we can integrate (exactly or approximately) the tau-delta equations and also the equations for the $F_{i}$, the remaining steps to achieving a full solution would then be to impose the conditions
$\psi|\psi>+<\phi|\phi>=N\ ;\ \ \ \ \ <\psi|\psi>-<\phi|\phi>=\tau\ ;$ (99)
and
$\phi|\psi>=\gamma.$ (100)
IV Conclusions and outlook
In a previous paper cf23 , we began an exploration of a particular form of non-linear extension of quantum mechanics. The extension could be applied to any underlying dynamical system, and featured the introduction of two state vectors instead of the usual single state vector in ordinary quantum mechanics. The model was simple, and the effects of the non-linearity could be determined completely, although the meaning of the second state vector remained uncertain. An analogy was noted with the two-state formalism of Aharonov and Vaidman Aharonov-Vaidman which is a time-symmetric formulation of ordinary quantum mechanics, based on earlier work by Aharonov, Bergmann and Lebowitz Aharonov64 . Because of our model’s simplicity, the dynamics it incorporated was limited. In particular, the effects of the non-linearity persisted for only a finite duration, decaying exponentially for large times. In addition, the natural extension of the Born rule to the model failed to encompass any of the non-linear behavior. Therefore, in this paper, we have constructed an expanded version, finding the most general such model consistent with three defining principles that were characteristic of the earlier work. Although we are not yet able to solve this new model completely, we can already see, by studying some simple examples, that the time dependence of the state vectors is much richer than that allowed by the earlier model. In particular, closed orbits exist for a range of parameters, which was not possible when the non-linearity only persisted for a finite time.
Our work so far suggests 2 approximations that may be useful in making further progress. In this paper, we showed that, for a suitably chosen range of parameters, the evolution of $\kappa=\ln|\gamma|^{2}$ is governed by a Hamiltonian whose potential,
$V(\kappa)=4e^{\kappa}+ce^{p\kappa}$ (101)
rises exponentially for $|\absolutevalue{\kappa}\rightarrow\infty$. We solved for the state vectors in the case that $\kappa$ remains stationary at the minimum of the potential. Building on this solution, we can study small oscillations of $\kappa$ about the minimum, to see how this affects the dynamics of the state vectors themselves, i.e. we set
$\kappa=\kappa_{0}+\delta\kappa$ (102)
where $e^{\kappa_{0}}=\left(\frac{-\mu}{4b}\right)N^{2}$ is the minimum of the potential, and work to lowest order in $\delta\kappa$.
Clearly, this approximation requires $\kappa$ to remain close to the bottom of the potential. A complementary approximation, which should be better the farther $\kappa$ gets from the bottom, is to replace the true potential $V(\kappa)$ with
$\widetilde{V}(\kappa)=4e^{\kappa}\theta\left(\kappa-\kappa_{1}\right)+ce^{p\kappa}\theta\left(\kappa_{1}-\kappa\right)$ (103)
Here $\kappa_{1}$ is defined by $4e^{\kappa_{1}}=ce^{p\kappa_{1}}$, so that $\widetilde{V}$ is continuous (this is not the same as the minimum of the potential, unless p=1).p=-1). The motivation for this approximation is, first, that in each of the two regions $\kappa>\kappa_{1}$ and $\kappa<\kappa_{1}$ the difference between $V(\kappa)$ and $\widetilde{V}(\kappa)$ is an exponential tail that grows smaller the larger $|\kappa|$ becomes; and second, in each region the problem is solvable, because the potential is given by a single exponential, which is just the case considered in our first paper. We anticipate implementing both these approximations in future work.
Our point of view is rather unconventional. Our first goal is to see whether the non-linearities of our model can account for some or all of the classical behavior of gravity. Whether or not that hope is realized, further analysis of the extension of quantum mechanics that we have proposed may yield valuable insight into the as yet unsolved problem of what quantum mechanics is really all about.